OpenStudy (anonymous):

In the reaction, CH4 + 2O2 arrow CO2 + 2H2O, 890 kilojoules of heat are released. Using Einstein's equation, calculate the mass amount to which this energy equates. (1 J = 1 kg m2 s-2) and the speed of light = 3.0 x 108 m/sec Answer 9.9 x 1012 kg 9.9 x 10-12 kg 9.9 x 10-14 kg 1.1 x 10-6 kg

OpenStudy (anonymous):

In the equation E = mc^2, what is m? In this case, that is.

OpenStudy (anonymous):

@abb0t

OpenStudy (anonymous):

@JFraser

OpenStudy (jfraser):

m is what you're looking for, since the problem says "what mass"

OpenStudy (jfraser):

E is given, 889kJ, and the speed of light is constant 3*10^8 m/s

OpenStudy (anonymous):

oh... m = E x c?

OpenStudy (jfraser):

no

OpenStudy (anonymous):

but where does the ^2 go?

OpenStudy (jfraser):

since \[E = mc^2\] then \[ m = \frac{E}{c^2}\]

OpenStudy (jfraser):

the squared function is part of the speed of light, you can't just ignore it when you rearrange an equation

OpenStudy (anonymous):

ah....right.. sorry, I haven't had food yet, my brain's running on low with a subject it already has a hard time with. so... m = 890/3 x 108^2?

OpenStudy (jfraser):

almost. the energy is given in kJ, but has to be in J for the equation. You actually have: \[m = \frac{8.89*10^5J}{(3.0*10^8\frac{m}{s})^2}\]

OpenStudy (jfraser):

you have to be careful where your exponents are when you divide in scientific notation

OpenStudy (anonymous):

hmm... I got 8.001 x 10^22, which isn't an option.. confusion?

OpenStudy (jfraser):

then your math is wrong, because you should get an incredibly small number

OpenStudy (jfraser):

you should review how to manipulate scientific notation and the rules of multiplying and dividing with exponents

OpenStudy (anonymous):

Oh, wait...no I think the answer is B

OpenStudy (jfraser):

do out the calculation and see

OpenStudy (anonymous):

yeah, it's B, thnx for your help!

OpenStudy (jfraser):

it is, GJ